# Qualification-free convex analysis via the joint supporting subspace

### v4

@@ 6 A framework for local face lattices
We now show the above statement. Let $C^{\prime}=C\cap(V+S)$, $\bar{C}=\operatorname{\Pi}_{U^{\perp}}C^{\prime}$, and $\bar{U}=\operatorname{\Pi}_{U^{\perp}}(U+S)$. Note that $\bar{C}$ is a convex set and $\bar{U}$ a singleton $\{x\}\subseteq\bar{C}$. We argue that $\bar{U}$ is a face of $\bar{C}$. Arguing by contradiction, we hypothesize the $\bar{U}$ is not a face of $\bar{C}$, in which case there are $\bar{y},\bar{z}\in\bar{C}\setminus\bar{U}$ such that $x=\lambda\bar{y}+(1-\lambda)\bar{z}$ for $\lambda\in(0,1)$. This implies that there are elements $y,z\in C^{\prime}$ such that $\bar{y}=\operatorname{\Pi}_{U^{\perp}}y$ and $\bar{z}=\operatorname{\Pi}_{U^{\perp}}z$, with $w=\lambda y+(1-\lambda)z$. Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$. This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Corollary 6.1, means that $U\notin\mathcal{F}(C^{\prime};S)$. This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show. We know that $U\in\mathcal{F}(C;S)$, so by Lemma 6.3, $U=U\cap V\in\mathcal{F}(C\cap(V+S);S)$, i.e., $U\in\mathcal{F}(C^{\prime};S)$. Therefore, we indeed have a contradiction, from which we find that $\bar{U}$ is a face of $\bar{C}$.
@@ 6.1 Minimal supporting subspace
The r.h.s. is the intersection of two convex faces of $C$. Their intersection is then also a face (this is immediate from Definition 3.1). By Theorem 6.1, $C\setminus[(U\cap V+S)\cap C]$ is convex. The convexity of this set, which we can write as $C\setminus((U\cap V)+S)$, implies that $U\cap V\in\mathcal{F}(C;S)$. ∎
@@
By Theorem 6.1, $H_{C}(S)$ being the smallest supporting subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own supporting subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a supporting subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is a supporting subspace, as implied by Theorem 6.1. Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other supporting subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$. ∎
@@
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated supporting subspace $H_{C}(S)$ is the only supporting subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
@@ 6.2 Nested normals of a difference of sets
In this section, we show the following important result.
Theorem 6.6 (Characterization of the nested normals of a difference of sets).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and
@@
We argue by structured induction over the supporting subspaces of CDC-D, as described in Corollary 6.2. Fixing some $S\subseteq C\cap D$, the statement $A:\mathcal{F}(C-D;0)\to\{\text{true},\text{false}\}$ that we show is the following: $A(U)$ is true if, and only if, all of the following hold.
  1. 1.
    $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S).$
  2. 2.
    $N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
  3. 3.
    $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$ The additional point 3) has a supportive role in this proof.
@@ 7.1 Regularity in the joint supporting subspace
We now show the main property of the joint supporting subspace, Theorem 4.2. We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [mordukhovich_VariationalAnalysisApplications_2018, Theorem 2.19].
@@ 7.2 Analytical characterizations
$=\operatorname{\mathrm{span}}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D}+C\cap D)$
$=\operatorname{aff}(F_{C}\cup F_{D}).$
@@ 7.4 Acknowledgements
The author is supported by the University of British Columbia. The author has no relevant financial or non-financial interests to disclose.

### v5

@@ 6 A framework for local face lattices
We now show the above statement. Let $C^{\prime}=C\cap(V+S)$, $\bar{C}=\operatorname{\Pi}_{U^{\perp}}C^{\prime}$, and $\bar{U}=\operatorname{\Pi}_{U^{\perp}}(U+S)$. Note that $\bar{C}$ is a convex set and $\bar{U}$ a singleton $\{x\}\subseteq\bar{C}$. We argue that $\bar{U}$ is a face of $\bar{C}$. Arguing by contradiction, we hypothesize the $\bar{U}$ is not a face of $\bar{C}$, in which case there are $\bar{y},\bar{z}\in\bar{C}\setminus\bar{U}$ such that $x=\lambda\bar{y}+(1-\lambda)\bar{z}$ for $\lambda\in(0,1)$. This implies that there are elements $y,z\in C^{\prime}$ such that $\bar{y}=\operatorname{\Pi}_{U^{\perp}}y$ and $\bar{z}=\operatorname{\Pi}_{U^{\perp}}z$, with $w=\lambda y+(1-\lambda)z$. Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$. This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Proposition 6.2, means that $U\notin\mathcal{F}(C^{\prime};S)$. This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show. We know that $U\in\mathcal{F}(C;S)$, so by Lemma 6.3, $U=U\cap V\in\mathcal{F}(C\cap(V+S);S)$, i.e., $U\in\mathcal{F}(C^{\prime};S)$. Therefore, we indeed have a contradiction, from which we find that $\bar{U}$ is a face of $\bar{C}$.
@@ 6.1 Minimal supporting subspace
The r.h.s. is the intersection of two convex faces of $C$. Their intersection is then also a face (immediate from the definition of faces). By Proposition 6.1, $C\setminus[(U\cap V+S)\cap C]$ is convex. The convexity of this set, which we can write as $C\setminus((U\cap V)+S)$, implies that $U\cap V\in\mathcal{F}(C;S)$. ∎
@@
By Proposition 6.1, $H_{C}(S)$ being the smallest supporting subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own supporting subspace (by Proposition 6.2), which would have to be properly contained in $H_{C}(S)$.
For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a supporting subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is a supporting subspace, as implied by Proposition 6.1. Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other supporting subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$. ∎
@@
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated supporting subspace $H_{C}(S)$ is the only element in $\mathcal{F}(C;S)$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
@@ 6.2 Nested normals of a difference of sets
We characterize the supporting subspaces of a difference of sets CDC-D from the supporting subspaces and nested normals of the individual sets $C$ and $D$ at $C\cap D$. One might hope for the simple statement that $\mathcal{F}(C;S)\cap\mathcal{F}(D;S)=\mathcal{F}(C-D;0)$ for $S\subseteq C\cap D$. Unfortunately, though $(\supseteq)$ turns out to hold, $(\subseteq)$ does not in general.
Example 6.1 (Mismatched nested normals).
Let $h:\mathbb{R}\to\mathbb{R},h(x):=x^{2}$, $C:=\operatorname{epi}\min(\iota_{\mathbb{R}_{-}},h)$ and $D:=\mathrm{hypo}\max(-\iota_{\mathbb{R}_{-}},-h)$, where $\mathrm{hypo}$ denotes the hypograph of the concave function. Then $\{0\}$ is a face of both $C$ and $D$, yet $C-D=\mathbb{R}\times\mathbb{R}_{+}$, and therefore $\{0\}\notin\mathcal{F}(C-D;0)$.
There is therefore a need for some additional condition on supporting subspaces shared by $C$ and $D$ for them to also be supporting subspaces of CDC-D. The next result allows us to derive such a condition by using nested normals.
Theorem 6.4 (Decomposition of supporting subspaces of differences).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then
  1. 1.
    $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S).$
  2. 2.
@@
  3. 3.
    $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$
The next corollary shows that the missing condition is the existence of a synchronized “path” of nested normals for $C$ and $D$.
Corollary 6.2 (Characterization of the faces of differences of convex sets).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C\cap D$ non-empty. The elements of $\mathcal{F}(C-D;0)$ are precisely the elements $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$ for which there exists a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$ such that $V_{0}=\mathbb{R}^{n}$, $V_{k}=U$, and such that for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\cap-N_{D}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
Proof of Corollary 6.2.
That any $U$ with such a sequence is in $\mathcal{F}(C-D;0)$ follows by first noticing that each $v_{i}$ in the sequence belongs to $N^{U}_{C-D}(0)$ by Theorem 6.4. By applying Theorem 6.1 sequentially, it follows that $U\in\mathcal{F}(C-D;0)$.
That any element $U\in\mathcal{F}(C-D;0)$ has such a sequence follows from Theorem 6.2, which shows that there is a sequence of nested normals of CDC-D at 0 that, starting from $\mathbb{R}^{n}$, attains $U$. By Theorem 6.4, this sequence corresponds to a sequence $v_{i},\ldots,v_{k}$ as described in the statement. ∎
We can now better understand Example 6.1. The supporting subspace $\{0\}$ of $C$ and $D$ is obtained via the sequence of nested normals $\{(0,-1),(1,0)\}$ for $C$ and $\{(0,1),(1,0)\}$ for $D$. Only the first nested normals are opposite to each other, and therefore there is no synchronized path of nested normals to $\{0\}$. If we modify the example by taking $D:=\mathrm{hypo}\max(-\iota_{\mathbb{R}_{+}},-h)$, then $\{0\}$ is attained by $\{(0,1),(-1,0)\}$, and the sequence is synchronized with that of $C$. In this modified example, $\{0\}$ is indeed a supporting subspace of CDC-D because CDC-D is a convex set such that $C-D\subseteq\mathbb{R}\times\mathbb{R}^{+}$ and $C-D\cap(\mathbb{R}\times\{0\})=\mathbb{R}_{-}\times\{0\}$, which shows that $\{0\}\in\mathcal{F}(C-D;0)$.
@@
We argue by nested normal induction (from Corollary 6.1) over the supporting subspaces of CDC-D. Fixing some $S\subseteq C\cap D$, let $A(U)$ be the statement that all three points in the statements hold for supporting subspace $U\in\mathcal{F}(C-D;0)$.
@@ 7.1 Regularity in the joint supporting subspace
We now show the main property of the joint supporting subspace, Theorem 4.2. We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [19, Theorem 2.19].
@@ 7.2 Analytical characterizations
$=\operatorname{\mathrm{span}}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D}+C\cap D)=\operatorname{aff}(F_{C}\cup F_{D}).$
@@ 7.3 Second characterization
7.4 Conclusion
In this work we define and characterize the joint supporting subspace $T_{a}$, which exists between any two convex sets. By defining this affine subspace from the domains of two convex functions, and then reducing each function to it, qualification conditions are guaranteed to hold. We obtain a number of characterizations for $T_{a}$ by bridging between the facial structure of the difference of two convex sets CDC-D and the facial structure of the individual convex sets. By doing this, we show that “pathology on the boundary” of convex functions can be tamed by considering the facial structure of their domains.
Future work avenues include specifying and implementing efficient reduction algorithms for convex functions, and pairing them with optimization solvers to obtain robust algorithms with strong theoretical guarantees.
@@ 7.5 Acknowledgements
The author is supported by the University of British Columbia.